# Dockformer: A transformer-based molecular docking paradigm for large-scale virtual screening

### v2

@@ Abstract
Molecular docking enables virtual screening of compound libraries to identify potential ligands that target proteins of interest, a crucial step in drug development; however, as the size of the compound library increases, the computational complexity of traditional docking models increases. Deep learning algorithms can provide data-driven research and development models to increase the speed of the docking process. Unfortunately, few models can achieve superior screening performance compared to that of traditional models. Therefore, a novel deep learning-based docking approach named Dockformer is introduced in this study. Dockformer leverages multimodal information to capture the geometric topology and structural knowledge of molecules and can directly generate binding conformations with the corresponding confidence measures in an end-to-end manner. The experimental results show that Dockformer achieves success rates of 90.53% and 82.71% on the PDBbind core set and PoseBusters benchmarks, respectively, and more than a 100-fold increase in the inference process speed, outperforming almost all state-of-the-art docking methods. In addition, the ability of Dockformer to identify the main protease inhibitors of coronaviruses is demonstrated in a real-world virtual screening scenario. Considering its high docking accuracy and screening efficiency, Dockformer can be regarded as a powerful and robust tool in the field of drug design.
@@
Traditional docking approaches use scoring functions to measure the binding affinity of a given protein–ligand complex and then find the best binding conformation by applying optimization algorithms to minimize these functions [5]. For example, GOLD uses a genetic algorithm to search complex conformations [6], and AutoDock combines a genetic algorithm with a simulated annealing algorithm [7]. Although these optimization-based docking methods are commonly used in modern drug designs because of their good usability and interpretability, they still face the following challenges: the scoring functions are generally not precise enough, and optimization algorithms cannot guarantee that the global optimum is found each time. Although several advanced methods can offer reliable binding affinity predictions [8, 9], docking approaches require multiple independent optimization processes to sample possible binding conformations for each protein–ligand pair, leading to very high computational costs in LSVS tasks [2, 10].
@@
Inspired by the groundbreaking advancement of AlphaFold2 in protein structure prediction [11], a series of deep learning (DL)-based methods have emerged to solve molecular docking tasks [12]. According to the types of neural network architectures, they can be divided into three categories: graph neural networks (GNNs)-based [13], transformer-based [14] and diffusion model-based docking methods [15]. The primary motivation of these studies is twofold: first, improving the ligand docking accuracy with the aid of the power learning capabilities of DL technologies, and second, speeding up the screening process by directly predicting ligand binding conformations to skip the time-consuming optimization of traditional docking approaches [16]. However, although tremendous efforts have been made to develop DL-based docking tools, few can perform well in docking accuracy and screening speed simultaneously due to inadequate generalizability and non-end-to-end architectures [17, 18]. Therefore, how to use DL models to generate protein-ligand binding conformations precisely and efficiently is still an open question in LSVS tasks.
@@
In this study, a novel transformer-based architecture named Dockformer, which is shown in Fig. 1, is proposed. Dockformer uses two separate encoders to leverage multimodal information to generate latent embeddings of proteins and ligands and can thus effectively capture molecular geometric details, including 2D graph topology and 3D structural knowledge. A binding module is then employed to detect intermolecular relationships effectively on the basis of learned latent embeddings. Finally, in the structure module, the established relationships are utilized to generate the complex conformations directly, and the coordinates of the ligand atoms are calculated in an end-to-end manner. In addition, the corresponding confidence measures of each generated conformation are utilized to distinguish binding strengths instead of traditional scoring functions. Dockformer achieves superior docking accuracy compared with that of state-of-the-art DL-based and optimization-based docking methods and simultaneously speeds up the conformation generation process by orders of magnitude. Thus, this method is capable of meeting the rapid throughput requirements of LSVS tasks. Dockformer, as a robust and reliable protein–ligand docking approach, may significantly reduce the development cycle and cost of drug design.
@@ II Related work
DL-based molecular docking methods can be divided into three main categories. The models in the first class encode proteins and ligands as graphs and use equivariant GNNs to predict intermolecular binding interactions [19]. For instance, DeepDock utilizes GNNs to construct a mixture density network, which is based on the distance likelihood of ligand-target node pairs and can act as a scoring function [20]. Then, DeepDock can accurately search complex conformations by optimizing the scoring function. EquiBind employs a SE(3)-equivariant GNN to detect the interactions between protein residues and ligand atoms, and uses gradient descent algorithms to determine the translation, rotation and torsion of binding conformations [13]. Similarly, TankBind uses a trigonometry-aware GNN to predict protein-ligand intermolecular distances. Then, it adopts a multi-dimensional scaling method to reconstruct the ligand atom coordinates based on the pair distances [21]. KarmaDock combines the methodologies of DeepDock and EquiBind, which utilizes a graph transformer neural network to learn pair distance distributions and employs E(n)-equivariant GNNs to generate binding conformations directly [22]. For molecular docking tasks, graph models can directly handle the structural geometry and effectively process the symmetry properties of molecular representations, enabling the movement direction and amplitude of ligand atoms to be updated in each message passing iteration. However, the oversmoothing issue results in the inadequate generalizability of these GNN-based methods. The performance of the methods has not yet reached that of conventional docking methods [17].
@@
The models in the second class are based on transformer architectures, which can efficiently capture long-range dependencies among intra- and intermolecular tokens. For example, Uni-Mol has pioneered the use of atom and pair representations to encode ligands and protein pockets, and employed the self-attention layers with pair biases to share information between each representation. It generates the coordinates of ligand atoms via two distance matrices predicted by pairwise representations [14]. Inspired by Alphafold2, GAABind incorporates the additional triangular self-attention layers into the main architecture of Uni-Mol, which can capture the geometric and topological properties of binding pockets and ligands [23]. In addition, considering transformer models typically exist urgent requirements of enormous training data, CarsiDock and HelixDock customized large-scale complex structure datasets for pretraining and used the crystallized structure dataset for fine-tuning to improve their generalization abilities [24, 25]. Although these transformer-based models can achieve satisfactory docking accuracy, most still require independent optimization procedures to generate binding poses from the predicted interaction distance maps or distributions. Therefore, their inference processes remain time-consuming and are expensive for LSVS tasks [18]. In addition, these models output physically implausible conformations with steric clashes and incorrect bond lengths and angles since they ignore the essential topological information of molecules during the conformation generation process [12].
@@
Unlike the aforementioned models, which treat molecular docking as regression problems, models in the third class frame docking problems as generative modeling tasks. Specifically, DiffDock uses a denoising diffusion probabilistic model over the non-Euclidean manifold of ligand conformations, and then maps the manifold to predict the translation, rotation and torsion of ligands [15]. DynamicBind adopts an equivariant geometric diffusion network to construct a smooth energy landscape, which can be used to recover ligand conformations based on the unbound structures of proteins [26]. Furthermore, NeuralPLexer employed a diffusion-based generative model to predict complex structures by solely inputting protein sequences and ligand molecular graphs [27]. AlphaFold3, which was recently proposed, applies diffusion transformer models as decoders to simultaneously calculate each atom coordinate of proteins and ligands and yields very remarkable prediction accuracy in molecular docking tasks [28]. However, such generative models require the sampling of many noisy conformations for prediction, leading to very high computational complexity and slow docking speed. Despite the superior prediction performance of AlphaFold3, its inference time is longer than that of most docking approaches, making it incapable of virtually screening the billions of compounds in large-scale libraries.
@@ III-A Architecture Overview
The essence of molecular docking lies in detecting nonbonded interactions between the atoms of a ligand and a protein pocket. The protein is conventionally regarded as a rigid body to simplify the calculations, and docking algorithms aim to generate the binding conformation of ligands within the protein–ligand complex on the basis of the predicted atom interactions. Consequently, the proposed Dockformer algorithm is designed to directly predict the 3D coordinates of all heavy atoms of the ligands for a given protein pocket. An architecture overview of Dockformer is depicted in Fig. 1. Dockformer consists of three stacked primary blocks. First, two independent encoders are used to encode the multimodal information of the ligand and binding pocket, and intramolecular interactions are exploited to produce their intrinsic representations. Second, a binding module captures intermolecular interactions between the binding pocket and ligand to generate the corresponding latent embeddings. Finally, the latent embeddings are fed into the structure module to predict the binding conformation of the ligand by considering the precise 3D coordinates of each atom.
@@ III-B Featurization Methodology
The network architecture simultaneously incorporates the 1D sequence, 2D graph and 3D geometry information of the ligand and protein pocket as inputs, enabling valuable insights from distinct modalities. Let denote the atom features used to encode the sequence information, where is the number of heavy atoms and represents the atom type of the -th atom by using a one-hot encoding scheme. Additionally, 2D graph information is encoded as the chemical bonds and structural interconnections between atom pairs in the ligand. Two-dimensional graph pair features contain two representations. First, represents the connection feature, which uses the shortest path distance between atoms and to reflect their connection relation in the graph. Second, records the edge feature to reflect the bond information. Denoting the edges along the shortest path of atoms and as , the edge feature can be calculated by , where are the learnable parameters. Notably, both and need to be calculated for the protein pocket, which is considered rigid during the docking process.
Next, the 3D geometric information of the ligand or protein pocket is encoded as two representations. Regarding the atoms as point clouds, the first representation is the learnable global position embedding to reflect spatial information according to the 3D coordinate of each atom, where and denote the dimensional number. of the -th atom can be calculated by
@@ III-C Encoder Modules
Two encoders are used to update the representations of both the ligand and protein pockets, which share the same architecture but have different weights. The architecture of the encoder module is depicted in Fig. 2. Specifically, the atom embeddings are initialized with the atom features and the global position embedding described by
@@ III-F Loss Functions
The training process of Dockformer is divided into two phases. First, the outputs of the binding module are mapped to predict the distances between atom pairs. The loss function of the predicted intra- and intermolecular distances against the corresponding ground truth distances can be quantified as follows:
where and denote the ground truth intramolecular and intermolecular distances, respectively. and are the atom numbers of the ligand and protein products, respectively, and and are the atom indices. An loss function is employed for minimizing the intramolecular distance error, whereas a robust loss function is used for minimizing the intermolecular error [30], presented as:
After the first training phase, the encoders and binding module can produce effective latent embeddings for both the ligand and protein pockets by considering the interactions between each other. Since the structure module can generate complex conformations in an end-to-end manner, the loss function with respect to the coordinates of the ligand atoms against the corresponding ground truth coordinates is incorporated during the second training phase, described by
where is the output coordinate of the structure module and where denotes the ground truth coordinate of the -th atom in the cocrystal structure. Because the structure module will be recycled times, only the coordinate loss of the last iteration is used to refine the complex conformation without extra computational cost.
In addition, considering that Dockformer was developed to screen small-molecule compounds for a specific target protein virtually, allocating confidence assessment indicators for each generated ligand–protein complex conformation will be constructive. Inspired by the confidence measure in AlphaFold2 [11], the distance difference test between the predicted distance and ground truth distance is used to calculate the target confidence of the predicted conformation.
where is obtained through the same projection head with atomic representations and pairwise representation in Eq. (11). The confidence indicator can be defined as follows:
@@
III-G Benchmark Datasets and Implementation Details
Similarly to most DL-based docking approaches, Dockformer is trained with the latest version of the well-established PDBbind V2020, which includes the cocrystal structures and the corresponding experimentally determined binding affinities of 19443 protein-ligand complexes released before 2020 [31]. The dataset is divided into training and validation sets with a partition ratio of 9:1, using the same filtering protocol provided in Uni-Mol [14]. In addition, the core set of PDBbind (also termed CASF-2016), which contains 285 hand-curated high-resolution complexes, is used to evaluate the docking ability of Dockformer. Dockformer is further evaluated on an independent test dataset named PoseBusters, which was recently developed and includes 428 crystal complexes released since 2021 [12]. By using this dataset, overlap between the training and test datasets is avoided, enhancing the rationality and reliability of the evaluations. In addition, for the proposed Dockformer, the pockets are used as Dockformer’s inputs instead of entire proteins to reduce the computational complexity because the binding sites of target proteins have been well studied in most virtual screening tasks. Even without exact pockets, some prediction algorithms can also effectively identify potential binding sites of proteins. Dockformer is trained using eight NVIDIA RTX A6000 GPUs and two 128-core Intel(R) Xeon(R) Gold 6338 CPUs @ 2.00 GHz. In both training phases, the adaptive moment estimation (Adam) optimizer is used to minimize the loss functions smoothly with a learning rate of , and an early stopping mechanism based on validation error is employed to prevent overfitting with predefined patience of 20 epochs.
@@ IV EXPERIMENTS
IV-A Prediction Performance on Benchmark Datasets

### v4

@@ Abstract
Molecular docking is a crucial step in drug development, which enables the virtual screening of compound libraries to identify potential ligands that target proteins of interest. However, the computational complexity of traditional docking models increases as the size of the compound library increases. Recently, deep learning algorithms can provide data-driven research and development models to increase the speed of the docking process. Unfortunately, few models can achieve superior screening performance compared to that of traditional models. Therefore, a novel deep learning-based docking approach named Dockformer is introduced in this study. Dockformer leverages multimodal information to capture the geometric topology and structural knowledge of molecules and can directly generate binding conformations with the corresponding confidence measures in an end-to-end manner. The experimental results show that Dockformer achieves success rates of 90.53% and 82.71% on the PDBbind core set and PoseBusters benchmarks, respectively, and more than a 100-fold increase in the inference process speed, outperforming almost all state-of-the-art docking methods. In addition, the ability of Dockformer to identify the main protease inhibitors of coronaviruses is demonstrated in a real-world virtual screening scenario. Considering its high docking accuracy and screening efficiency, Dockformer can be regarded as a powerful and robust tool in the field of drug design.
@@
Traditional docking approaches use scoring functions to measure the binding affinity of a given protein–ligand complex and then find the best binding conformation by applying optimization algorithms to minimize these functions [5]. For example, GOLD uses a genetic algorithm to search complex conformations [6], and AutoDock combines a genetic algorithm with a simulated annealing algorithm [7]. Although these optimization-based docking methods are commonly used in modern drug designs because of their good usability and interpretability, they still face the following challenges: the scoring functions are generally not precise enough, and optimization algorithms cannot guarantee that the global optimum is found every time. Although several advanced methods can offer reliable binding affinity predictions [8, 9], docking approaches require multiple independent optimization processes to sample possible binding conformations for each protein–ligand pair, leading to very high computational costs in LSVS tasks [2, 10].
@@
Inspired by the groundbreaking advancement of AlphaFold2 in protein structure prediction [11], a series of deep learning (DL)-based methods have emerged to solve molecular docking tasks [12]. These approaches can be divided into three categories according to their neural network architectures: graph neural networks (GNNs)-based [13], transformer-based [14] and diffusion model-based docking methods [15]. The primary motivation of these studies is twofold: first, improving the ligand docking accuracy with the aid of the powerful learning capabilities of DL technologies, and second, speeding up the screening process by directly predicting ligand binding conformations to skip the time-consuming optimization of traditional docking approaches [16]. Although tremendous efforts have been made to develop DL-based docking tools, few can perform well in docking accuracy and screening speed simultaneously due to inadequate generalizability and non-end-to-end architectures [17, 18]. In addition, existing methods solely focus on 1D sequential, 2D graph topological or 3D structural in isolation, and fail to leverage the integration and complementarity of each modality for capturing the inherent interactions between proteins and ligands. Therefore, how to use DL models to generate protein-ligand binding conformations precisely and efficiently is still an open question in LSVS tasks.
@@
In this study, a novel transformer-based architecture named Dockformer is proposed to overcome the above-mentioned issues of current DL-based docking methods. Specifically, Dockformer uses two separate encoders to leverage multimodal information to generate latent embeddings of proteins and ligands and can thus effectively capture molecular geometric details, including 2D graph topology and 3D structural knowledge. A binding module is then employed to detect intermolecular relationships effectively on the basis of learned latent embeddings. Finally, in the structure module, the established relationships are utilized to generate the complex conformations directly, and the coordinates of the ligand atoms are calculated in an end-to-end manner. In addition, the corresponding confidence measures of each generated conformation are utilized to distinguish binding strengths instead of traditional scoring functions. In summary, distinct from conventional DL-based and optimization-based docking methods, the multimodal information fusion equips Dockformer with superior docking accuracy, and the end-to-end architecture enables it to simultaneously speed up the conformation generation process by orders of magnitude. Thus, this method can meet the rapid throughput requirements of LSVS tasks. Dockformer, as a robust and reliable protein-ligand docking approach, may significantly reduce the development cycle and cost of drug design.
@@ II Related work
DL-based molecular docking methods can be divided into three main categories: GNNs-based, transformer-based and diffusion-based methods. The models in the first class encode proteins and ligands as graphs and use equivariant GNNs to predict intermolecular binding interactions [19]. For instance, DeepDock utilizes GNNs to construct a mixture density network, which is based on the distance likelihood of ligand-target node pairs and can act as a scoring function [20]. Then, DeepDock can accurately search complex conformations by optimizing the scoring function. EquiBind employs a SE(3)-equivariant GNN to detect the interactions between protein residues and ligand atoms, and uses gradient descent algorithms to determine the translation, rotation and torsion of binding conformations [13]. Similarly, TankBind uses a trigonometry-aware GNN to predict protein-ligand intermolecular distances. Then, it adopts a multi-dimensional scaling method to reconstruct the ligand atom coordinates based on the pair distances [21]. KarmaDock combines the methodologies of DeepDock and EquiBind, which utilizes a graph transformer neural network to learn pair distance distributions and employs E(n)-equivariant GNNs to generate binding conformations directly [22]. For molecular docking tasks, graph models can directly handle the structural geometry and effectively process the symmetry properties of molecular representations, enabling the movement direction and amplitude of ligand atoms to be updated in each message passing iteration. However, the over-smoothing issue results in the inadequate generalizability of these GNN-based methods. The performance of the methods has not yet reached that of conventional docking methods [17].
@@
The models in the second class are based on transformer architectures, which can efficiently capture long-range dependencies among intra- and intermolecular tokens. For example, Uni-Mol has pioneered the use of atom and pair representations to encode ligands and protein pockets, and employed the self-attention layers with pair biases to share information between each representation. It generates the coordinates of ligand atoms via two distance matrices predicted by pairwise representations [14]. Inspired by Alphafold2, GAABind incorporates the additional triangular self-attention layers into the main architecture of Uni-Mol, which can capture the geometric and topological properties of binding pockets and ligands [23]. In addition, considering transformer models typically exist urgent requirements of enormous training data, CarsiDock and HelixDock customized large-scale complex structure datasets for pretraining and used the crystallized structure dataset for fine-tuning to improve their generalization abilities [24, 25]. Although these transformer-based models can achieve satisfactory docking accuracy, most still require independent optimization procedures to generate binding poses from the predicted interaction distance maps or distributions. Therefore, their inference processes remain time-consuming and are expensive for LSVS tasks [18]. Furthermore, these transformer-based methods do not adequately account for the positional embeddings of individual tokens (atoms), which ultimately compromises the model’s generalizability and overall performance. In addition, these models output physically implausible conformations with steric clashes and incorrect bond lengths and angles since they ignore the essential topological information of molecules during the conformation generation process [12].
@@
Unlike the aforementioned models, which treat molecular docking as regression problems, models in the third class frame docking problems as generative modeling tasks. Specifically, DiffDock uses a denoising diffusion probabilistic model over the non-Euclidean manifold of ligand conformations, and then maps the manifold to predict the translation, rotation and torsion of ligands [15]. DynamicBind adopts an equivariant geometric diffusion network to construct a smooth energy landscape, which can be used to recover ligand conformations based on the unbound structures of proteins [26]. Furthermore, NeuralPLexer employed a diffusion-based generative model to predict complex structures by solely inputting protein sequences and ligand molecular graphs [27]. AlphaFold3, which is recently proposed, applies diffusion transformer models as decoders to simultaneously calculate each atom coordinate of proteins and ligands and yields very remarkable prediction accuracy in molecular docking tasks [28]. However, such generative models require the sampling of many noisy conformations for denoising step by step, leading to very high computational complexity and slow docking speed. Despite the superior prediction performance of AlphaFold3, its inference time is longer than that of most docking approaches, making it incapable of virtually screening the billions of compounds in large-scale libraries.
To overcome the drawbacks of conventional DL-based docking algorithms, Dockformer first uses two separate encoders to integrate multimodal information for generating latent embeddings of proteins and ligands. Each encoder effectively captures molecular geometric details from 2D graph topology and 3D structural knowledge, enabling a more comprehensive understanding of molecular interactions. It is because 2D graph information allows us to grasp the bonding relationships and connectivity patterns that are crucial for accurate docking predictions, and the 3D structural information provides spatial context, ensuring an account for the actual conformational geometry of molecules. Second, Dockformer uses an end-to-end decoder to generate the complex conformations and the corresponding confidence measures directly. It can skip time-consuming optimization and denoising processes, thereby significantly accelerating the docking procedure and enhancing computational efficiency. The confidence measures can be used to distinguish binding strengths, replacing traditional scoring functions that may not accurately reflect true binding affinities. These properties empower Dockformer to achieve superior docking accuracy and screening efficiency in the LSVS tasks, compared with state-of-the-art DL-based algorithms.
@@ III-A Architecture Overview
The essence of molecular docking lies in detecting nonbonded interactions between the atoms of a ligand and a protein pocket. The protein is conventionally regarded as a rigid body to simplify the calculations, and docking algorithms aim to generate the binding conformation of ligands within the protein–ligand complex on the basis of the predicted atom interactions. Consequently, the proposed Dockformer algorithm is designed to directly predict the 3D coordinates of all heavy atoms of the ligands for a given protein pocket. The architecture of the proposed Dockformer is depicted in Fig. 1. Dockformer consists of three stacked primary blocks. First, two independent encoders are used to encode the multimodal information of the ligand and binding pocket, and intramolecular interactions are exploited to produce their intrinsic representations. Second, a binding module captures intermolecular interactions between the binding pocket and ligand to generate the corresponding latent embeddings. Finally, the latent embeddings are fed into the structure module to predict the binding conformation of the ligand by considering the precise 3D coordinates of each atom.
@@ III-B Featurization Methodology
The network architecture simultaneously incorporates the 1D sequence, 2D graph and 3D geometry information of the ligand and protein pocket as inputs, enabling valuable insights from distinct modalities. Let denote the initial atom features used to encode the sequence information, where is the number of heavy atoms and represents the atom type of the -th atom by using a one-hot encoding scheme. Additionally, 2D graph information is encoded as the chemical bonds and structural interconnections between atom pairs in the ligand. Two-dimensional graph pair features contain two representations. First, represents the connection feature, which uses the shortest path distance between atoms and to reflect their connection relation in the graph. Second, records the edge feature to reflect the bond information. Denoting the edges along the shortest path of atoms and as , the edge feature can be calculated by , where are the learnable parameters. Notably, both and need to be calculated for the protein pocket, which is considered rigid during the docking process.
Next, the 3D geometric information of the ligand and protein pocket is encoded as two representations. Regarding the atoms as point clouds, the first representation is the learnable global position embedding to reflect spatial information according to the 3D coordinate of each atom, where and denotes the dimensional number. of the -th atom can be calculated by
@@ III-C Encoder Modules
Two encoders are used to update the representations of both the ligand and protein pockets, which share the same architecture but have different weights. The architecture of the encoder module is depicted in Fig. 2. Specifically, the atom embeddings of the first layer are initialized with the atom features and the global position embedding , as described by
@@ III-E Structure Module
This end-to-end framework enables Dockformer to directly learn the inherent mapping from input molecular structures to the desired docking conformations. That avoids the need for redundant iterative gradient descent optimization and denoising processes typically required in traditional DL-based docking approaches, significantly reducing computational overhead and greatly saving docking time.
@@ III-F Loss Functions
The training process of Dockformer is divided into two phases. First, the outputs of the binding module are mapped to predict the distances between atom pairs. The loss function of the predicted intra- and intermolecular distances against the corresponding ground-truth distances can be quantified as follows:
where and denote the ground-truth intramolecular and intermolecular distances, respectively. and are the atom numbers of the ligand and protein products, respectively. and are the atom indices. loss function is employed for minimizing the intramolecular distance error, whereas a robust loss function is used for minimizing the intermolecular error [30], presented as:
After the first training phase, encoders and binding module can produce effective latent embeddings for both the ligand and protein pockets by considering the interactions between each other. Since the structure module can generate complex conformations in an end-to-end manner, the loss function with respect to the coordinates of the ligand atoms against the corresponding ground-truth coordinates is incorporated during the second training phase, described by
where is the output coordinate of the structure module and where denotes the ground-truth coordinate of the -th atom in the cocrystal structure.
In addition, considering that Dockformer is developed to screen small-molecule compounds for a specific target protein virtually, allocating confidence assessment indicators for each generated ligand–protein complex conformation will be constructive. Inspired by the confidence measure in AlphaFold2 [11], the distance difference test between the predicted distance and ground-truth distance is used to calculate the target confidence of the predicted conformation.
where is obtained through the same projection head with atomic representations , denotes the different thresholds and pairwise representation in Eq. (11). The confidence indicator can be defined as follows:
@@ IV EXPERIMENTS
Extensive experiments are conducted to thoroughly evaluate the effectiveness of Dockformer. Specifically, Section IV-A describes the experimental setup, detailing the data preparation, model configurations and evaluation metrics. In Section IV-B, the docking accuracy of Dockformer is compared against state-of-the-art methods on Benchmarks. Section IV-C explores the computational complexity of Dockformer to highlight how the end-to-end framework accelerates the docking process. In Section IV-D, the generated confidence measures are assessed to showcase its ability to reliably distinguish binding strengths, without relying on traditional scoring functions. Section IV-E demonstrates Dockformer’s applicability in an LSVS task to emphasize its scalability and efficiency in real-world drug discovery scenarios. Section IV-F discusses the implementation of physical plausibility correction, ensuring that the generated conformations are physically realistic and chemically valid. Finally, Section IV-G presents the ablation experiments to analyze the contributions of each component to the overall performance and generalization ability of Dockformer, including multimodal fusion and structure module.
IV-A Experimental Setup
Similarly to most DL-based docking approaches, Dockformer is trained with the latest version of the well-established PDBbind V2020, which includes the cocrystal structures and the corresponding experimentally determined binding affinities of 19443 protein-ligand complexes released before 2020 [31]. The dataset is divided into training and validation sets with a partition ratio of 9:1, using the same filtering protocol provided in Uni-Mol [14]. In addition, the core set of PDBbind (also termed CASF-2016), which contains 285 hand-curated high-resolution complexes, is used to evaluate the docking ability of Dockformer. Dockformer is further evaluated on an independent test dataset named PoseBusters, which is recently developed and includes 428 crystal complexes released since 2021 [12]. By using this dataset, overlap between the training and test datasets is avoided, enhancing the rationality and reliability of the evaluations. In addition, for the proposed Dockformer, the pockets are used as Dockformer’s inputs instead of entire proteins to reduce the computational complexity because the binding sites of target proteins have been well studied in most virtual screening tasks. Even without exact pockets, some prediction algorithms can also effectively identify potential binding sites of proteins. Dockformer is trained using eight NVIDIA RTX A6000 GPUs and two 128-core Intel(R) Xeon(R) Gold 6338 CPUs @ 2.00 GHz. In both training phases, the adaptive moment estimation (Adam) optimizer is used to minimize the loss functions smoothly with a learning rate of , and an early stopping mechanism based on validation error is employed to prevent overfitting with predefined patience of 20 epochs. In addition, to ensure the fairness of experiments, the configurations of each optimization-based algorithm are maintained according to the default settings, and the parameters of DL-based methods are strictly configured as their original implementation details. That can reduce potential biases introduced by inconsistent parameter tuning and ensure the validity of the performance evaluation results.
IV-B Prediction Performance on Benchmark Datasets